# GNN-based Anchor Embedding for Efficient Exact Subgraph Matching

### v4

@@ Abstract
Index Terms:
Subgraph matching, graph isomorphism, graph processing, matching optimization.
@@ II Problem Definition
A graph $G=(V,E,L,\Sigma)$ is isomorphic to a graph $G^{\prime}=(V^{\prime},E^{\prime},L^{\prime},{\Sigma}’)$ ’ ), denoted as $G\cong G^{\prime}$, if there exists a bijecton $f:V\to V^{\prime}$ such that (1) $\forall v\in V$, $L(v)=L^{\prime}(f(v))$; (2) $\forall u,v\in V$, $(u,v)\in E$ if and only if $(f(u),f(v))\in E^{\prime}$.
@@ III An Overview
/* build indexes over graph $G$/
@@ IV-B GNN Model for Anchored Subgraph Embedding
GNN Model Architecture. Fig. 5 illustrates the GNN model. Specifically, we adopt a classic Graph Isomorphism Network (GIN) [16]) model for subgraph matching. Here are the reasons: (1) For the above condition C1, the GNN model (e.g., GIN) naturally has the order-invariant property [13] to enable isomorphic anchored subgraphs to have the same embedding. We prove this in Theorem IV.1. (2) For the above condition C2, the GIN model is as strong as the Weisfeiler-Lehman (WL) graph isomorphism testing to make non-isomorphic anchored subgraphs have different embeddings as much as possible, and stronger than other embedding methods, such as the graph factorization embedding method [17]. Theorem IV.2 provides a rigorous theoretical guarantee. (3) The GIN model has a simple architecture and high execution efficiency. Therefore, the classic GIN model is suitable for subgraph matching. If a more advanced GNN model satisfies conditions C1 and C2, and has higher execution efficiency than GIN in the future, the GIN model can be replaced by it.
@@
A linear layer performs a linear transformation for $h^{(t)}_{G}\in\mathbb{R}^{n\times 1}$ ($n$ is the $readout$ hidden dimension) and obtains the embedding vector $o(S^{k}_{G}(u,v))$ with dimension $m$ for $S^{k}_{G}(u,v)$.
@@
GNN Model Training. For each edge $(u,v)$ with $d(u)\leq d^{*}\lor d(v)\leq d^{*}$, we extract all its possible anchored subgraphs. In real-world graphs, vertex degrees usually follow the power-law distribution [19], and only a small fraction of vertices have high degrees [11]. For example, 85% of vertices have degree $\leq 10$ in DBLP graph data [1]. Therefore, it is usually acceptable for these edges with low vertex degrees to extract all possible anchored subgraphs.
@@
[…]
@@
The proof of Lemma IV.1 can be found in [16]. It shows that classic GNN can be as powerful as the WL test in distinguishing non-isomorphic graphs. Thus, we have Theorem IV.1.
@@
If $G\cong G^{\prime}$, they must satisfy the WL test [20, 16]. Then, the proof immediately follows from the inverse negative proposition of Lemma IV.1 that if the WL test decides that $G\cong G^{\prime}$, a graph neural network $\mathcal{M}$ maps $G$ and $G^{\prime}$ to the same embedding vector (i.e., $o(G)=o(G^{\prime})$). ∎
@@
The GNN Training w.r.t. The Embedding Quality: We train a GNN model (with 2 GIN layers modeled by a 1-layer MLP, SUM as the readout function, the $readout$ hidden dimension $n=10$, and the output embedding vector dimension $m=3$) over a large training data set $D$, where by default $|V(G)|=500K$, $avg\_deg(G)=5$ and the vertex label domain size $|\Sigma|=100$. The learning rate of the Adam optimizer $\eta=0.001$ and the batch size is $1K\sim 2K$.
@@
Fig. 7 reports the anchored 1-radius subgraph embedding conflict ratio of our GNN-AE. We vary $avg\_deg(G)$ from 3 to 10 in Fig. 7(a) and adjust $|V(G)|$ from $5K$ to $1M$ in Fig. 7(b). We adjust the training epochs $\rho\in\{5,10,20\}$. From the Fig. 7, the anchored subgraph embedding conflict ratio via the GNN model is not greater than $10^{-5}\%$. This means that almost any two non-isomorphic anchored subgraphs are mapped to non-conflicting embeddings.
Complexity Analysis of the GNN Training. For each GIN layer, let the number of layers in the MLP be $\beta$. The initial feature embedding dimension of the input and the hidden dimension in the MLP are both $n$. Assume we obtain the anchored 1-radius subgraphs in the data graph $G$. For an anchored subgraph $S$ in $D$, $|E(S)|\leq d_{v}$ and $|V(S)|=|E(S)|+1$. The time complexity of the computation on the GIN layer is $O(\beta\cdot n^{2}\cdot|V(S)|+|E(S)|)=O(\beta\cdot n^{2}\cdot d_{v})$. The upper bound of the number of anchored subgraphs in the training data set $D$ is $\sum_{v\in V(G)\land d_{v}\leq d^{*}}{(2^{d_{v}}-1)}$. The time complexity of the $readout$ function is $O(n\cdot|V(S)|)$, and the linear layer is $O(m\cdot n)$. We set the number of training epochs to be $\rho$. Thus, the total time complexity of the GNN training is $O(\rho\cdot\sum_{v\in V(G)\land d_{v}\leq d^{*}}{2^{d_{v}}}\cdot(\beta\cdot n^{2}\cdot d_{v}+m\cdot n))$.
@@ V-C Matching Growth Algorithm
In MatchGrowth procedure, the time complexity is given by $O(\sum_{(q_{i},q_{j})\in E_{dfs}(Q)}{(|C_{q_{i}q_{j}}^{S}|+|C_{q_{i}q_{j}}^{S^{\prime}}|+|C_{q_{i}q_{j}}^{P}|)})$, where $|C_{q_{i}q_{j}}^{P}|$ is the number of dense-dense candidate matches. Since GNNs have a high embedding quality with rigorous theoretical guarantees, the subgraph embedding conflict ratio via the GNNs is extremely low. For example, the conflict ratio is less than $10^{-5}\%$ on both real and synthetic data graphs in experiments. This means that almost all non-isomorphic subgraphs have unique embeddings without conflicts. As a result, each query edge can obtain a compact candidate set (i.e., fewer invalid results) that matches feature subgraphs. Thus, fewer invalid intermediate results are generated in the matching growth procedure compared to existing methods.
@@ VII-A Experimental Setup
Baseline Methods. We compare the performance of GNN-AE with 7 representative subgraph matching baselines: GraphQL (GQL) [23], QuickSI (QSI) [24], RI [25], Hybrid (mixed by CFL [26], RI [25] and GQL) [1], RapidMatch (RM) [7], BICE [27], and GNN-PE [11]. The RL-QVO [12] method relies on historical query graphs to train the model to work. The model must be retrained when queries change, which limits its robustness and scalability. Our GNN-AE does not rely on historical queries to work (our model training only relies on data graphs), which has a natural advantage over RL-QVO. Thus, we do not take RL-QVO as one of the baselines.
@@
$avg\_deg(G)$ of synthetic data graph $G$ 3, 4, 5, 6, 8
@@
Query Graphs. We follow previous works [7, 1, 26, 28, 11, 27] to generate query sizes and types for each data graph $G$. Specifically, for each dataset, we perform a random walk on $G$ until getting the specified number of vertices and extract the induced subgraph. Table III lists the size $|V(Q)|$ of query $Q$. Except for queries with $|V(Q)|=4$, each query set on real-world graphs contains two categories: dense query graph (i.e., $avg\_deg(Q)>3$) and sparse query graph (i.e., $avg\_deg(Q)\leq 3$). For each query size and category, we generate 100 queries.
@@ VII-B Efficiency and Effectiveness Evaluation
In subsequent experiments, we also evaluate the efficiency of our GNN-AE under varying query and data graph characteristics (e.g., query graph patterns, $|V(Q)|$, $|V(G)|$, $avg\_deg(G)$, and $|\Sigma|$ of $G$). To better show the curve trends, we omit the baseline results in subsequent experiments.
Online Computation Cost: Fig. 16 reports the computation cost of four steps (including DFS query plan, obtain feature embeddings, get candidate matches, and match growth) in the online subgraph matching phase (i.e., Algorithm 3). We select 80 query graphs of ye and wn respectively, and record the elapsed time of four steps. We find that the get candidate matches and match growth steps take up most of the elapsed time for subgraph matching. For the DFS query plan and obtaining feature embeddings, the time cost of our GNN-AE remains low (i.e., $<1.7$ $ms$, and $<6.5$ $ms$, respectively).
The GNN-AE Efficiency w.r.t Query Graph Patterns. Fig. 17 presents the total query time of our GNN-AE on different query graph patterns on wn dataset, including dense (i.e., $avg\_deg(Q)>3$, denoted as $D$) and sparse (i.e., $avg\_deg(Q)\leq 3$, denoted as $S$) query graphs with varying sizes. The figure shows that our GNN-AE can perform better than baselines on most query graph patterns. Especially, for most query patterns, GNN-AE can perform better than traditional exploration-based backtracking search methods (e.g., GQL, QSI, and Hybrid) and BICE by 1-2 orders of magnitude.
@@
The GNN-AE Scalability w.r.t Data Graph. Fig. 19 evaluates the total query time of GNN-AE for all queries with different data graph characteristics, including the data graph size $|V(G)|$, $avg\_deg(G)$, and $|\Sigma|$ of $G$. Intuitively, larger $|V(G)|$ and $avg\_deg(G)$, more candidate matches for the query, which incurs an increased time cost. When $|\Sigma|$ increases, the feature embeddings have a stronger filtering power, which produces fewer candidate matches for the query. Thus, the elapsed time of GNN-AE shows a downward trend. Nevertheless, for varying data graphs, GNN-AE still remains a low time cost (i.e., $<47$ $sec$), which confirms the scalability of GNN-AE.

### v5

@@ II Problem Definition
[…]
@@ III An Overview
/* build indexes over graph $G$ */
@@ IV-A Feature Subgraphs
In real-world graphs, vertex degrees usually follow the power-law distribution [16], and only a small fraction of vertices have high degrees [11]. For example, 85% of vertices have degree $\leq 10$ in DBLP graph data [1]. Therefore, we can set $d^{*}=10$. Specifically, when the maximum vertex degree $d_{max}(u)$ of the data graph is large (e.g., $d_{max}(u)>8$), we adopt the anchored 1-radius subgraph to avoid the combinatorial explosion of the number of anchored substructures. According to Lemma IV.1, the edge $(u,v)$ has $2^{d(u)-1}$ possible anchored 1-radius subgraphs. We can further directly deduce that the edge $(u,v)$ with $d(u)\leq d^{*}\lor d(v)\leq d^{*}$ has at most $2^{d^{*}-1}$ anchored 1-radius subgraphs. When $d^{*}=10$, there are no more than 512 ($2^{9}$) anchored subgraphs for edge $(u,v)$, which is acceptable. When the maximum vertex degree $d_{max}(u)$ of the data graph is not large (e.g., $d_{max}(u)\leq 8$), we can adopt the anchored 2-radius or 3-radius subgraph to obtain a better filtering effect in subgraph matching due to their more complex structures. For such data graphs, it is also acceptable to extract all possible anchored 2-radius or 3-radius subgraphs. In fact, there are lots of data graphs/networks with a low maximum vertex degree in the real world, such as DNA molecular graph structures [17], urban road networks [18], IoT sensor networks in a smart park [19], base station networks for cellular mobile communications [20, 21], and so on.
•  DNA molecular graph structure: A DNA molecule contains millions of nucleotides, each of which can serve as a vertex in the graph structure. Each nucleotide is linked to the nucleotide above and below it by a phosphodiester bond, and to the nucleotide in the complementary chain through a base. Therefore, the maximum degree of each nucleotide vertex is only 3.
•  Urban road network: A road can be considered as an edge in the graph, and a road intersection as a vertex. In the road network, there are many types of intersections, such as crossroads, three-way intersections, and highway ramps. Each road intersection type can be mapped to a unique vertex label. In the real world, the number of forks at a road intersection usually does not exceed 8. Thus, the road network is a typical graph structure with a low maximum vertex degree.
•  IoT sensor network in a smart park: Vertices are sensors in the Internet of Things (IoT), such as temperature sensors, security cameras, and smart parking detectors. The number of vertices in an IoT sensor network of a smart park can reach hundreds of thousands. Due to the limitations of wireless transmission distance and energy consumption, a sensor is usually only connected to 3-5 adjacent sensors for data backup or relay transmission. Therefore, the maximum degree of an IoT sensor network in a smart park is usually only 5.
•  Base station network for cellular mobile communications: Each base station is a vertex. To avoid signal interference and ensure coverage, a base station usually only establishes data connections with 5-6 neighboring base stations. Thus, the base station network is another a typical data graph with a low maximum vertex degree.
Lemma IV.1.
For an edge $(u,v)$ in graph $G$, if vertex $u$ has degree $d(u)$, there are $2^{d(u)-1}$ possible anchored 1-radius subgraphs.
Proof.
According to Definition IV.1, the anchored 1-radius subgraph with respect to $(u,v)$ is a subgraph of $G$ that must contain the edge $(u,v)$ and some vertices $w$ with distance $d(u,w)\leq 1$. Let the 1-hop subgraph starting from vertex $u$ be $S=(V_{S},E_{S},L_{S},\Sigma_{S})$. The number of possible anchored 1-radius subgraphs of $(u,v)$ is equivalent to the number of subsets of the non-$(u,v)$ edge sets $\Theta=\{e|e\in E_{S}\wedge e\neq(u,v)\}$ in $S$, where $|\Theta|=d(u)-1$. Thus, there are $C^{0}_{d(u)-1}+C^{1}_{d(u)-1}+\dots+C^{d(u)-1}_{d(u)-1}=2^{d(u)-1}$ possible anchored 1-radius subgraphs. The lemma holds. ∎
@@ IV-B GNN Model for Anchored Subgraph Embedding
GNN Model Architecture. Fig. 5 illustrates the GNN model. Specifically, we adopt a classic Graph Isomorphism Network (GIN) [23]) model for subgraph matching. Here are the reasons: (1) For the above condition C1, the GNN model (e.g., GIN) naturally has the order-invariant property [13] to enable isomorphic anchored subgraphs to have the same embedding. Theorem IV.1 also provides a rigorous proof. Some existing graph encoding or embedding methods cannot satisfy the condition C1, such as DeepWalk [24], and GraphSAGE [26]. Because they introduce the randomization step (e.g., sampling, random walks) into their execution. (2) For the above condition C2, the GIN model is as strong as the Weisfeiler-Lehman (WL) graph isomorphism testing to make non-isomorphic anchored subgraphs have different embeddings as much as possible, and stronger than other embedding methods, such as the graph factorization embedding method [27]. Theorem IV.2 provides a rigorous theoretical guarantee. (3) The GIN model has a simple architecture and high execution efficiency. Therefore, the classic GIN model is suitable for subgraph matching. If a more advanced GNN model satisfies conditions C1 and C2, and has higher execution efficiency than GIN in the future, the GIN model can be replaced by it.
@@
Property IV.1 indicates that no matter how the vertices in the graph $G$ are numbered, the output of the function $f$ is the same. That is, the output of function $f$ regardless of the assignment of vertex IDs on the graph $G$. The GNN model satisfies the order-invariant property, which means that the embedding generated by the GNN focuses on the topological structure of the graph $G$ rather than the vertex ID assignment. This is the theoretical basis for the GNN model to support graph isomorphism testing.
@@
A linear layer performs a linear transformation for $h^{(t)}_{G}\in\mathbb{R}^{n\times 1}$ ($n$ is the readoutreadout hidden dimension) and obtains the embedding vector $o(S^{k}_{G}(u,v))$ with dimension $m$ for $S^{k}_{G}(u,v)$.
@@
GNN Model Training. For each edge $(u,v)$ with $d(u)\leq d^{*}\lor d(v)\leq d^{*}$, we extract all its possible anchored subgraphs. As mentioned in subsection IV-A, when the maximum vertex degree $d_{max}(u)$ of the data graph is large (e.g., $d_{max}(u)>8$), we adopt the anchored 1-radius subgraph to avoid the combinatorial explosion of the number of anchored substructures. When the maximum vertex degree $d_{max}(u)$ of the data graph is not large (e.g., $d_{max}(u)\leq 8$), we can adopt the anchored 2-radius or 3-radius subgraph to obtain a better filtering effect. For such data graphs, it is also acceptable to extract all possible anchored 2/3-radius subgraphs.
@@
Training Data Set: Algorithm 2 describes the GNN model training for anchored subgraph embedding. As described in subsection IV-A, all possible anchored subgraphs for each edge $(u,v)$ with $d(u)\leq d^{*}\lor d(v)\leq d^{*}$ in the data graph $G$ constitute the training data set $D$. In order to make non-isomorphic anchored subgraphs learn different embeddings as much as possible via the GNN model, we assign a unique $m$-dimensional ($m\geq 1$) vector as a label to each anchored subgraph in $D$, and make the vectors far enough from each other (e.g., the distance at least $\xi$). If $m=1$, when $D$ is large, the 1-dimensional vector results in the label range being too wide, which is not conducive to model convergence. For example, in Fig. 6(a), in order to keep the distance between vectors at least $\xi$, the 1-dimensional labels have a wide range. However, when $m>1$, labels can be restricted to a small space (e.g., a rectangle or a cube in Fig. 6(b)(c).), and kept far enough from each other, which is beneficial for model convergence. Thus, we adopt an $m$-dimensional vector as the label, where $m>1$. Specifically, we randomly choose a range $[a,b]$, and generate $\lceil\sqrt[m]{|D|}\rceil$ values $\{r_{1},r_{2},\dots,r_{\lceil\sqrt[m]{|D|}\rceil}\}$ from $[a,b]$, where the distance between any two values is at least $\xi$. The value of each dimension in the $m$-dimensional vector is $r\in\{r_{1},r_{2},\dots,r_{\lceil\sqrt[m]{|D|}\rceil}\}$. Therefore, we can combine ${\lceil\sqrt[m]{|D|}\rceil}^{m}\geq|D|$ types of $m$-dimensional vectors, and the Euclidean distance between any two vectors is at least $\xi$. These vectors are assigned to anchored subgraphs in $D$ as labels.
@@
The proof of Lemma IV.2 can be found in [23]. It shows that classic GNN can be as powerful as the WL test in distinguishing non-isomorphic graphs. Thus, we have Theorem IV.1.
@@
If $G\cong G^{\prime}$, they must satisfy the WL test [29, 23]. Then, the proof immediately follows from the inverse negative proposition of Lemma IV.2 that if the WL test decides that $G\cong G^{\prime}$, a graph neural network $\mathcal{M}$ maps $G$ and $G^{\prime}$ to the same embedding vector (i.e., $o(G)=o(G^{\prime})$). ∎
@@
The GNN Training w.r.t. The Embedding Quality: We train a GNN model (with 2 GIN layers modeled by a 1-layer MLP, SUM as the readout function, the readoutreadout hidden dimension $n=10$, and the output embedding vector dimension $m=3$) over a large training data set $D$, where by default $|V(G)|=500K$, avg_deg(G)=5avg\_deg(G)=5 and the vertex label domain size $|\Sigma|=100$. The learning rate of the Adam optimizer $\eta=0.001$ and the batch size is $1K\sim 2K$.
@@
Fig. 7 reports the anchored 1-radius subgraph embedding conflict ratio of our GNN-AE. We vary avg_deg(G)avg\_deg(G) from 3 to 10 in Fig. 7(a) and adjust $|V(G)|$ from $5K$ to $1M$ in Fig. 7(b). We adjust the training epochs $\rho\in\{5,10,20\}$. From the Fig. 7, the anchored subgraph embedding conflict ratio via the GNN model is not greater than $10^{-5}\%$. This means that almost any two non-isomorphic anchored subgraphs are mapped to non-conflicting embeddings.
Complexity Analysis of the GNN Training. For each GIN layer, let the number of layers in the MLP be $\beta$. The initial feature embedding dimension of the input and the hidden dimension in the MLP are both $n$. Assume we obtain the anchored 1-radius subgraphs in the data graph $G$. For an anchored subgraph $S$ in $D$, $|E(S)|\leq d_{v}$ and $|V(S)|=|E(S)|+1$. The time complexity of the computation on the GIN layer is $O(\beta\cdot n^{2}\cdot|V(S)|+|E(S)|)=O(\beta\cdot n^{2}\cdot d_{v})$. The upper bound of the number of anchored subgraphs in the training data set $D$ is $\sum_{v\in V(G)\land d_{v}\leq d^{*}}{(2^{d_{v}}-1)}$. The time complexity of the readoutreadout function is $O(n\cdot|V(S)|)$, and the linear layer is $O(m\cdot n)$. We set the number of training epochs to be $\rho$. Thus, the total time complexity of the GNN training is $O(\rho\cdot\sum_{v\in V(G)\land d_{v}\leq d^{*}}{2^{d_{v}}}\cdot(\beta\cdot n^{2}\cdot d_{v}+m\cdot n))$.
@@ V-C Matching Growth Algorithm
In MatchGrowth procedure, the time complexity is given by $O(\sum_{(q_{i},q_{j})\in E_{dfs}(Q)}{(|C_{q_{i}q_{j}}^{S}|+|C_{q_{i}q_{j}}^{S^{\prime}}|+|C_{q_{i}q_{j}}^{P}|)})$, where $|C_{q_{i}q_{j}}^{P}|$ is the number of dense-dense candidate matches. Since GNNs have a high embedding quality with rigorous theoretical guarantees, the subgraph embedding conflict ratio via the GNNs is extremely low. For example, in Fig. 7, the conflict ratio is less than $10^{-5}\%$ on data graphs in experiments. This means that almost all non-isomorphic subgraphs have unique embeddings without conflicts. As a result, each query edge can obtain a compact candidate set (i.e., fewer invalid results) that matches feature subgraphs. Thus, fewer invalid intermediate results are generated in the matching growth procedure compared to existing methods.
@@ VII-A Experimental Setup
Baseline Methods. We compare the performance of our GNN-AE approach with 7 representative exact subgraph matching baseline methods: GraphQL (GQL) [32], QuickSI (QSI) [33], RI [34], Hybrid (mixed by CFL [35], RI [34] and GQL) [1], RapidMatch (RM) [7], BICE [36], and GNN-PE [11].
@@
avg_deg(G)avg\_deg(G) of synthetic data graph $G$ 3, 4, 5, 6, 8
@@
Query Graphs. We follow previous works [7, 1, 35, 37, 11, 36] to generate query sizes and types for each data graph $G$. Specifically, for each dataset, we perform a random walk on $G$ until getting the specified number of vertices and extract the induced subgraph. Table III lists the size $|V(Q)|$ of query $Q$. Except for queries with $|V(Q)|=4$, each query set on real-world graphs contains two categories: dense query graph (i.e., avg_deg(Q)>3avg\_deg(Q)>3) and sparse query graph (i.e., $avg\_deg(Q)\leq 3$). For each query size and category, we generate 100 queries.
@@ VII-B Efficiency and Effectiveness Evaluation
In subsequent experiments, we also evaluate the efficiency of our GNN-AE under varying query and data graph characteristics (e.g., query graph patterns, $|V(Q)|$, $|V(G)|$, avg_deg(G)avg\_deg(G), and $|\Sigma|$ of $G$). To better show the curve trends, we omit the baseline results in subsequent experiments.
Online Computation Cost: Fig. 16 reports the computation cost of four steps (including DFS query plan, obtain feature embeddings, get candidate matches, and match growth) in the online subgraph matching phase (i.e., Algorithm 3). We select 80 query graphs of ye and wn respectively, and record the elapsed time of four steps. We find that the get candidate matches and match growth steps take up most of the elapsed time for subgraph matching. For the DFS query plan and obtaining feature embeddings, the time cost of our GNN-AE remains low (i.e., <1.7<1.7 msms, and <6.5<6.5 msms, respectively).
The GNN-AE Efficiency w.r.t Query Graph Patterns. Fig. 17 presents the total query time of our GNN-AE on different query graph patterns on wn dataset, including dense (i.e., avg_deg(Q)>3avg\_deg(Q)>3, denoted as $D$) and sparse (i.e., $avg\_deg(Q)\leq 3$, denoted as $S$) query graphs with varying sizes. The figure shows that our GNN-AE can perform better than baselines on most query graph patterns. Especially, for most query patterns, GNN-AE can perform better than traditional exploration-based backtracking search methods (e.g., GQL, QSI, and Hybrid) and BICE by 1-2 orders of magnitude.
@@
The GNN-AE Scalability w.r.t Data Graph. Fig. 19 evaluates the total query time of GNN-AE for all queries with different data graph characteristics, including the data graph size $|V(G)|$, avg_deg(G)avg\_deg(G), and $|\Sigma|$ of $G$. Intuitively, larger $|V(G)|$ and avg_deg(G)avg\_deg(G), more candidate matches for the query, which incurs an increased time cost. When $|\Sigma|$ increases, the feature embeddings have a stronger filtering power, which produces fewer candidate matches for the query. Thus, the elapsed time of GNN-AE shows a downward trend. Nevertheless, for varying data graphs, GNN-AE still remains a low time cost (i.e., <47<47 secsec), which confirms the scalability of GNN-AE.